OpenStudy (anonymous):

What value of m solves the equation 2^m = 1/8

OpenStudy (anonymous):

Here, you must be knowing that: \[8 = 2 \times 2 \times 2 \implies 8 = 2^3\]

OpenStudy (anonymous):

So: \[\frac{1}{8} = \frac{1}{2^3} \implies \frac{1}{8} = 2^{-3}\]

OpenStudy (anonymous):

So, just equate the two things: \[\large 2^m = 2^{-3}\] Equate the exponents of left and right hand side to get m..

OpenStudy (anonymous):

@ari2403 I need your response here.. :)

OpenStudy (anonymous):

thank you so much